### v1

# Local Conditional Controlling for Text-to-Image Diffusion Models

Abstract
Diffusion models have exhibited impressive prowess in the text-to-image task. Recent methods add image-level controls, e.g., edge and depth maps, to manipulate the generation process together with text prompts to obtain desired images. This controlling process is globally operated on the entire image, which limits the flexibility of control regions. In this paper, we introduce a new simple yet practical task setting: local control. It focuses on controlling specific local areas according to user-defined image conditions, where the rest areas are only conditioned by the original text prompt. This manner allows the users to flexibly control the image generation in a fine-grained way. However, it is non-trivial to achieve this goal. The naive manner of directly adding local conditions may lead to the local control dominance problem. To mitigate this problem, we propose a training-free method that leverages the updates of noised latents and parameters in the cross-attention map during the denosing process to promote concept generation in non-control areas. Moreover, we use feature mask constraints to mitigate the degradation of synthesized image quality caused by information differences inside and outside the local control area. Extensive experiments demonstrate that our method can synthesize high-quality images to the prompt under local control conditions. Code is available at https://github.com/YibooZhao/Local-Control.
Figure 1: Previous image-level control and the proposed local control. Local control takes text prompt, image conditions and user-specified local control regions as input, generating images that are faithful to both the prompt and local control conditions.
1 Introduction
Large-scale text2image diffusion models [29, 27, 26] have showed remarkable capabilities in generating images conditioned on specified text prompts. However, text descriptions frequently fall short in efficiently or sufficiently conveying detailed control [38] over the final image generation results. Recent methods [44, 21] have introduced the incorporation of image-level spatial controls, such as edges, depth, and segmentation, into the text2image generation process. These methods operate across the entire image, ensuring that the final synthesized image retains the complete global structural information of the original. Such advancements have attracted considerable attention from both the academic and industrial sectors, as it facilitates a more controllable image synthesis process.
Unfortunately, the conventional approach of enforcing control at the entire-image level lacks the fine-grained image manipulation capability that users often require. Considering the example in Figure 1, we might want to exclusively apply the cat canny condition. This naturally leads to a question: Can we confine our control to specific local areas using image conditions? In this paper, we specifically address this practical scenario and introduce the concept of local control as a novel and pragmatic paradigm. It focuses on controlling the generation of specific areas of an image according to user-defined local control conditions, while the rest of the image is conditioned by the original text prompt. As illustrated in Figure 1, only the canny condition of the cat is employed for controlling, while the remainder of the image is generated solely based on the provided text prompt. This novel manner empowers users with the ability to exert fine-grained and flexible control over the image generation process.
An intuitive resolution for local control lies in directly adding local image conditions. However, this naive method leads to a catastrophic problem. The prompt concept that is most related to the local control condition dominates the generation process, while other prompt concepts are ignored. Consequently, the generated image cannot align with the input prompt. We call this problem local control dominance. We show the cases in the left column of Figure 2. Delving into the underlying mechanism, we find that the attention maps show high response areas for both the cat and dog within the local control regions, which impedes the generation of multiple objects, as seen here with the dog’s generation being hindered. It indicates that local control conditions force the model to focus on the controlled area and overlook other areas.
To address this issue, we introduce an object regeneration strategy. It identifies the ignored objects in the cross attention map, enhancing them on the latent by adaptively operating attention scores under local and non-local regions. This manner guides the model to generate the most related objects to the image condition in the local region, meanwhile producing other desired objects in non-control regions for the text prompt. Additionally, we suppress attention scores for patches on the cross attention map which lacks the strongest response in the token dimension. It aims to facilitate the generation of objects in non-local region. Moreover, we use local region mask to constraint the feature output of Controlnet. This technology alleviate the degradation of synthesized image quality caused by differences in control information between non local and local control regions. All proposed strategies are operated at the inference stage. Overall, the proposed training-free approach achieves high-quality local control for both the text prompt and image conditions, without the need for any training or additional text-image paired datasets. Extensive experiments demonstrate the effectiveness of our method.
Our contributions can be summarized as follows:
  • We introduce a new task setting: local control, allowing for fine-grained controlling with image conditions. It aims to control specific regions within the image as desired by the user, while the remaining areas are conditioned solely by the original text prompt.
  • We pose the main challenge in local control, and propose a training-free method to resolve it. The proposed approach ensures the preservation of local control while maintaining the original generative capability of the diffusion models.
  • Extensive experiments demonstrate that our method can synthesize high-quality images, precisely aligning both local image conditions and text prompts.
Figure 2: Local control dominance. On the left, we exhibit a case of local control dominance, illustrated by the ControlNet cross-attention maps for “cat” and “dog”. The response for ”dog” in non-local regions is lower, leading to its absence in the generated image. On the right, we display the results of our method applied to local control with the same seed, where both “cat” and “dog” are successfully generated.
2 Related work
2.1 Text-to-Image Models
Early methods [41, 43] for text-to-image synthesis are limited in generating low-resolution images in specific domains. Extensive data volume[31] and abundant model capacity have led to significant progress in the field of text-to-image synthesis. In recent years, the diffusion model [3, 12] has achieved amazing results in text-to-image generation tasks and has gradually become a representative method in this field. Such as Imagen [29], Stable Diffusion [27], and DALL-E [26]. They leverage the denoising process, gradual reduction of noise to high-quality images through iterative optimization. In addition, the cross-attention mechanism incorporates content from text prompts into the generated images through CLIP[24] or large language models[25]. However, they are not sensitive to every concept in text prompts. To mitigate this problem, a few works like StructureDiffusion [4], Attend-and-Excite [2] and Multi-Concept T2I-Zero [36] refine the cross-attention map to strengthen all of the subjects to encourage the model to generate all concepts described in the text prompts.
2.2 Compositional Generation
Image composition involves placing a foreground object onto a background image. Simply placing a foreground object onto a background can lead to an unconvincing result. However, more recent works [42, 1, 20] utilize diffusion models to tackle the problems of fusion inconsistency, enhancing the efficiency of image composition. [1] takes an input image and a mask, and adjusts the area covered by the mask to match a guiding text prompt while leaving the unmasked regions unchanged. TF-ICON[20] is introduced as a pipeline that employs readily available diffusion models for cross-domain image-guided composition, eliminating the need for extra training, fine-tuning, or optimization. Paint by Example [42] trains a diffusion model to merge a reference image onto a source image at a location indicated by a mask of any shape, utilizing classifier-free guidance to assist the process. Recent advancements include VisorGPT[40], LayoutGPT [5], LLM blueprint[9], and ControlGPT[45], which are models specifically designed for planning visual layouts in image synthesis.
2.3 Controllable Generation
To facilitate controllable image generation, many works focus on customization concepts[33, 14, 8, 15], multi-concepts[17, 39, 30, 10], or conditional control[44, 21]. Text-driven control methods focus on adjusting prompts and manipulating token embedding. For instance, textual-inversion [7] injects the concept into the text embedding to control specific subject generation. MakeAScene [6] utilizes semantic masks to guide image generation. DreamBooth [28], Perfusion [34] and Custom Diffusion [16] allow for the personalization of content in the generated image through fine-tuning the image diffusion model using a small set of user-provided example images. Prompt2Prompt [11] and Plug-and-Play [35] explore the relationship between attention map and image representation. These methods can refine attention maps for image editing purposes. While they achieve high synthesis quality, text prompts may not always provide reliable structural guidance for the synthesis results. PITI [38] proposes to enhance structural guidance by reducing the gap between features derived from conditions. [37] suggests using the similarity gradient between the target sketch and intermediate model features to impose constraints on the structure of the final results. As concurrent works, [44] learns condition-specific ControlNet to enable guide generation for the pre-trained diffusion model. T2i-Adapter[21] introduces a simple and lightweight adapter to combine the internal knowledge in T2I models with external control signals. We find that existing controllable image generation methods typically focus on image-level control, which lack the ability to specify local control for generating images that align with both prompts and control conditions. Addressing this gap, we introduce local control as a novel and pragmatic paradigm. It enables precise control over specific image areas according to user-defined conditions, while the remaining parts are guided by the original text prompts. Additionally, we propose a seamless, training-free method compatible with existing ControlNet frameworks. Under the paradigm of local control, this method allows for high-quality image generation that closely aligns with the descriptions in the prompt.
Figure 3: Overview of our method. Given the text condition and local control condition, a latent variable $z_{T}$ is passed into the denoising network, i.e., UNet. At each denoising step, we apply Feature Mask Constraint (see Section 3.6) to the ControlNet branch output features. Cross-attention maps generated from U-net are refined by Focused Token Response (see Section 3.5), which suppresses weaker interfering responses to facilitate the concept generation of non-local regions. Furthermore, we use Object Regeneration (see Section 3.4) to update the resulting latent $z_{t}$ using attention score differences $l$, to identify and regenerate ignored objects in the cross-attention map.
3 Method
3.1 Preliminaries
In this work, we apply our method on Stable Diffusion [27], which efficiently operates on the latent space of a pre-trained variational autoencoder. Specifically, given an input image $I$, the encoder $\mathcal{E}$ maps it to a latent $z$. image diffusion algorithms progressively add noise to the image and produce a noisy image $z_{t}$, where $t$ represents the number of times noise is added. Stable Diffusion enables the control of synthesized image content through additional textual inputs. The text prompt is first pre-processed to text tokens by the text encoder of pre-trained CLIP [22]. Diffusion model are trained using a squared error loss to learn a network $\epsilon_{\theta}$ that predicts the noise added to the noisy image $\bm{z}_{t}$:
$\mathcal{L}=\mathbb{E}_{\bm{z}_{0},\bm{t},\bm{c}_{t},,\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{c}_{t}))\|_{2}^{2}\Big{]}$ (1)
Relying solely on text prompts often proves insufficient for generating images that align with the user’s specific desires, recent methods integrate image-level spatial controls. ControlNet[44] utilizes the deep and robust encoding layers of a large-scale diffusion model. It employs zero convolutions that gradually increase the parameters from zero, ensuring that no detrimental noise impacts the fine-tuning process. During training, an additional spatial control condition $\bm{c}_{\text{f}}$ is required, and the training loss is:
$\mathcal{L}=\mathbb{E}_{\bm{z}_{0},\bm{t},\bm{c}_{t},\bm{c}_{\text{f}},\epsilon\sim\mathcal{N}(0,1)}\Big{[}\|\epsilon-\epsilon_{\theta}(\bm{z}_{t},\bm{t},\bm{c}_{t},\bm{c}_{\text{f}}))\|_{2}^{2}\Big{]}$ (2)
3.2 Overview
In the context of local control, the spatial structure information required for the local control region is designated by the user as $\mathcal{C}_{f}$. Given a text prompt, the tokenizer produces a set of tokens $\mathcal{P}=\{\mathbf{p}_{i}\}$, and a set of corresponding spatial cross-attention maps $\mathcal{A}^{t}=\{\mathbf{A}^{t}_{i}\}$ in Unet can be accordingly obtained. The most straightforward method of enabling local control is to directly input $\mathcal{C}_{f}$ and $\mathcal{P}$ into the existing control methods. However, this approach leads to the issue of local control dominance (Figure 2).
In this paper, we propose a training-free method that synthesizes images using local control condition and the prompt as inputs. This method builds upon prior research [11] which has shown that the prominently responsive regions within the cross-attention map are crucial in determining the spatial positioning of generated objects. Our method is illustrated in Figure 3. Specifically, our method consists of three components: 1) Object Regeneration (see Section 3.4), which facilitates the generation of ignored objects; 2) Focused Token Response (see Section 3.5), which suppresses non-maximum attention scores for removing interference to facilitate the concept generation in non-local region; 3) Feature Mask Constraint (see Section 3.6), which mitigates image quality degradation caused by discrepancies in control information.
3.3 Control Concept Matching
Our initial objective in the image synthesis process with local control is to identify the most suitable object for generation within the control region at timestep t. The resulting object token indices are denoted as $\text{C}^{t}_{\text{control}}$. In our method, the sum of attention scores within the local control region is employed as the criterion. At denoising steps $t>\beta T$, we identify the object with the highest sum attention score within the local control region as the $\text{C}^{t}_{\text{control}}$. $\beta$ is a hyperparameter that acts on the total timesteps $T$. Given the impact of the denoising process’s initial steps on the spatial structure of the synthesized image, we focus on stabilizing the selection of $\text{C}^{t}_{\text{control}}$. Thus, we employ the $Count_{max}(\cdot)$ function to determine the most frequently occurring $\text{C}^{t}_{\text{control}}$ in the early steps of the denoising process.
$\text{C}^{t}_{\text{control}}=\begin{cases}{\arg\max}_{i}\Big{(}\sum(\mathbf{M}\cdot\mathbf{A}^{t}_{i})\Big{)},&t>\beta T,\\ Count_{max}(\{\text{C}^{t^{*}}_{\text{control}}\mid t^{*}>\beta T\}),&t\leq\beta T\end{cases}$ (3)
where $\mathbf{M}$ denotes a binary spatial mask, which is generated from the local control region. It aims to selectively local control elements in the cross-attention maps. This approach ensures that the chosen object is optimally aligned with the user-defined control conditions.
Figure 4: Comparisons to ControlNet [44] with local image conditions.
3.4 Object Regeneration
Due to the impact of local control dominance, most objects often respond to the local control regions in cross-attention maps. An effective strategy to counter this effect involves directing the attention responses of objects in non-control regions away from the local control region. To achieve this, we introduce the object regeneration strategy, which is specifically designed to enhance the difference in response between the inside and outside of the local control region. We select a set of object tokens $\mathcal{S}=\{\mathbf{s}_{i}\}$ from the collection $\mathcal{P}$, where $\mathcal{P}$ represents the entire set of prompt tokens.
$\begin{split}L_{s_{i}}=D\big{[}&\max(G(\mathbf{A}_{s_{i}})\cdot\mathbf{M})\\ &-\max(G(\mathbf{A}_{s_{i}})\cdot(1-\mathbf{M}))\big{]}\end{split}$ (4)
When $s_{i}$ is the token corresponding to $\text{C}^{t}_{\text{control}}$, $D$ is set to -1; otherwise, it is 1. $G(\cdot)$ represents a Gaussian smoothing operator. $L_{s_{i}}$ uses the difference between the maximum attention scores inside and outside the local control region to represent distance.
$z_{t}=z_{t}-\alpha_{t}\nabla l$ (5)
$l$ is determined by $\max(L_{si})$ and is used to update the noised latent $z_{t}$ at the current time step $t$, where $\alpha_{t}$ is a scalar defining the step size of the gradient update. $l$ updates noised latents by increasing the attention scores that exhibit the maximum difference between patchs inside and outside the local control region. This approach directs the model to generate objects closely related to the control condition in the local region, while also creating objects in the non-control areas as given text prompt.
3.5 Focused Token Response
After updating the noised latent, while the image content closely aligns with the prompt, objects still tend to be generated near the local control region. This challenge is often linked to the overlapping contribution areas of corresponding token embeddings. To enhance object distinction, inspired by[36], we utilize cross attention maps to mitigate weak response regions. Specifically, we reduce the attention scores for patchs in the attention map that do not demonstrate the strongest response in the token dimension. This technology encourages the reduction of responses from less dominant tokens, thereby enforce the model to selectively respond to the most relevant tokens.
$\mathcal{A}^{t}=O(\arg\max(\mathcal{A}^{t}),\gamma)\cdot\mathcal{A}^{t}$ (6)
Here, $\arg\max$ identifies the indices of the highest responses across token dimension, while $O$ represents a suppression operation that applies a scaling factor $\gamma$ to the non-max attention scores. This process reduces the impact of irrelevant pixels in the cross attention map for the current token.
Figure 5: Visual results of our method under the local control setting. The proposed method can be employed for flexibly controlling the local regions with image conditions.
3.6 Feature Mask Constraint
During the denoising process with local control, there is a notable difference in the amount of information between inside and outside of the local control region. The local control carries only localized control information, leaving the rest of the areas with blank control information. As a consequence, the latent distribution resulting from the injection of local control conditions into the U-Net differs from that obtained under global control conditions, potentially affecting the overall image quality.
To address this issue, we apply feature mask constraint to the output of ControlNet. This design specifically directs the U-Net’s focus to the local control region when adding control. It thus minimizes the disruptive effect of irrelevant blank control information in non-control region. In practical terms, we extract the local feature maps from the output of ControlNet block. Subsequently, these local condition features were added in the Unet. Suppose $\mathcal{F}(\cdot)$ is a neural block in U-Net where $x$ is the input feature map. Separately, $\mathcal{G}(\cdot)$ is a neural block in ControlNet, and the conditional input feature $c_{f}$ is derived from $\mathcal{C}_{f}$. The resultant feature map $y$ includes added control conditions.
$y=\mathcal{F}(x)+\mathbf{M}\cdot\mathcal{G}(x,c_{f})$ (7)
After applying the feature mask constraint, the output from the ControlNet is selectively integrated into the U-Net. The local control condition is achieved through a multiplication operation involving the feature mask $\mathbf{M}$ and the conditioned output from the ControlNet block, $\mathcal{G}(x,c_{f})$. This combination is then added to the original feature map $\mathcal{F}(x)$ from the U-Net. By explicitly preventing the U-Net from focusing on blank control information, this method effectively mitigates image quality issues arising from disparities in control information.
4 Experiments
4.1 Dataset and Evaluation
Dataset: We utilized the COCO[19] validation set, which comprises 80 object categories. Each image in this set is associated with multiple captions; for our evaluation, we randomly selected one caption per image, resulting in a dataset of 5k generated images. Additionally, we create a dataset inspired by the Attend-and-Excite [2], named Attend-Condition. This dataset encompasses 11 object and 11 animal categories, with text prompts typically structured as ‘a [object] and a [animal]’. We selected one representative image for each concept to serve as control information for local control, generating two images per prompt, totaling 968 validation data.
Evaluation Metrics: For assessing image quality, we used the Fréchet Inception Distance (FID)[32], while the average CLIP[23] score between text prompts and corresponding images measured text-image similarity. The primary focus of our evaluation is on whether the generated images align with the objects mentioned in the captions. Utilizing the descriptive capabilities of BLIP-2[18], we generate captions for the images produced under the local control paradigm. Subsequently, we calculate the CLIP text-text similarity between these generated captions and the original captions, providing a comprehensive measure of content alignment with the text prompt.
Dataset Condition FID $\downarrow$ CLIP Score $\uparrow$ CLIP T2T $\uparrow$
ControlNet T2I-Adapter Ours ControlNet T2I-Adapter Ours ControlNet T2I-Adapter Ours
COCO2017 Canny 23.65 28.86 21.38 0.283 0.295 0.303 0.800 0.777 0.795
Depth 32.08 26.57 22.01 0.290 0.298 0.307 0.763 0.788 0.804
Sketch 25.96 34.06 22.33 0.303 0.292 0.307 0.799 0.771 0.806
Attend-Condition Canny - - - 0.271 0.254 0.287 0.750 0.694 0.802
Depth - - - 0.251 0.261 0.283 0.713 0.714 0.806
Sketch - - - 0.264 0.256 0.286 0.741 0.685 0.801
Table 1: Quantitative comparisons to other controllable methods: ControlNet [44] and T2I-Adapter [21]. We do not provide FID on the Attend-Condition dataset due to the small scale of its original images. Our method achieves the best performance under most metrics.
Figure 6: Comparisons to different controllable diffusion models on local control.
4.2 Visualization Results
Figure 4 demonstrates the practical utility of our method in achieving local control. It ensures that the generation process remains attentive to local content control, while simultaneously allowing the rest of the regions to generate content accurately in accordance with the prompt. Compared to images generated by ControlNet[44], our approach effectively reduces the issue of local control dominance.
Figure 5 further showcases the flexibility of our approach. It presents images generated under various conditions, showing the method’s ability to seamlessly integrate with the existing ControlNet framework. This flexibility highlights our approach’s capability to synthesize images accurately under a wide range of control conditions.
Figure 7: Ablation Studies. We use ControlNet[44] as our baseline. In the fourth column, we add Object Regeneration (OR) to baseline method. The fifth column builds upon the previous one by incorporating Focused Token Response (FTR). Finally, the sixth column extends the previous setup by adding Feature Mask Constraint (FMC).
4.3 Comparisons
We conducted a comparative analysis of our method against ControlNet[44] and T2I-Adapter[21], using local control condition as the input condition. This involved evaluating the methods under various conditions such as canny, depth, and sketch on the COCO and Attend-Condition datasets. The results, encompassing metrics FID, CLIP-score and CLIP text-text similarity (CLIP T2T), are detailed in Table 1. Due to the small number of original images in the Attend-Condition, we do not calculate its FID, as such a small sample size leads to unreliable FID. Our method outperformed the aforementioned control methods in most conditions under the paradigm of local control.
Figure 6 illustrates these differences, particularly in scenarios involving multiple objects in prompts. In the first example, ControlNet and T2I-Adapter tended to merge the concepts of a building and an airplane into a single object within the local control region. In contrast, our method generate a building in line with the local control conditions while accurately generating the plane in the non-local control area. Another striking example is seen in the third row on the left, demonstrates ControlNet’s output where it mistakenly generates part of a rabbit’s shape on a turtle. Our method, however, showcased its advanced capability to distinctly recognize and generate separate objects. This allows for greater flexibility in producing user-specified images within the local control.
4.4 Ablation Studies
In our ablation study, we utilize the canny condition on the COCO dataset to evaluate the effectiveness of our method’s key components. Specifically, we focus on three compon ents: 1) Object Regeneration, updates the noised latent to assist object generation in non-local regions; 2) Focused Token Response, designed to refine attention response; and 3) Feature Mask Constraint, a mechanism to constraint features of the ControlNet output. The results are listed in Table 2. It reveals that updating noised latent and refine attention response effectively reduce the influence of local control dominance in non-local regions. This enables more accurate image generation in accordance with the given prompts, as evidenced by a high CLIP T2T score of 0.806. Since non-local regions have blank control information, the gap in control conditions between inside and outside the local control regions impacts image quality. Our study shows that incorporating feature mask constraint significantly mitigates this issue. This integration resulted in notable improvements in both CLIP and FID scores, with CLIP scores rising to 0.303 and FID scores decreasing to 21.38, indicating enhanced overall image quality.
OR FTR FMC FID CLIP Score CLIP T2T
- - - 23.65 0.283 0.750
$\surd$ - - 24.48 0.299 0.786
$\surd$ $\surd$ - 22.90 0.298 0.806
$\surd$ $\surd$ $\surd$ 21.38 0.303 0.795
Table 2: Ablation studies on the proposed components, including Object Regeneration (OR), Focused Token Response (FTR), and Feature Mask Constraint (FMC). All components are crucial.
Figure 7 visualizes the impact of progressively integrating our designs on image generation. In the first row, using only the local control condition results in generating a train that matches its spatial structure, but the car is missing. Adding Attention Contrast Enhancement improves this, allowing the car mentioned in the prompt to appear, although some overlap and sticking issues are still evident. When Focused Token Response is incorporated, the image correctly produces all concepts from the prompt. Moreover, the introduction of the feature mask constraint significantly enhances image quality. Similar improvements are observed in other examples, confirming the effectiveness of our method’s components.
5 Limitations
The limitation pertains to the semantic connection between inside and outside of the local control region. This can cause problems with how objects are sized and placed, making them look out of proportion. Future work could address this by establishing a more comprehensive connection between the inside and outside of the local control regions.
6 Conclusions
In this paper, we introduce local control as a new paradigm for image controllable generation, focusing on user-defined control of specific image areas. To mitigate local control dominance, we propose a training-free method that integrates smoothly with existing control models. This approach optimizes noised latents and cross-attention maps, enhancing object generation in non-local regions. We demonstrate a variety of results under the local control condition, successfully generating high-quality images faithful to the text prompts.
Supplementary Material
Considering the space limitation of the main text, we provided more results and discussion in this supplementary material, which is organized as follows:
  • $\bullet$
    Section 7: implementation details.
  • $\bullet$
    Section 8: more ablation studies
    • discussion on Control Concept Matching.
    • selection of the $\gamma$ value in Focused Token Response
  • $\bullet$
    Section 9: more visualization results.
    • visualization results for COCO and Attend-Condition.
    • visualization results for different condition and prompt.
  • $\bullet$
    Section 10: the necessity of the local control.
7 Implementation Details
Experiments.
All experimental results are obtained using the official Stable Diffusion [27] v1.5 text-to-image model and ControlNet [44]. The number of denoising steps $T$ is set as 50 while configuring the classifier-free guidance to 7.5. $\beta$ is set as 0.8 in Control Concept Matching , which means that $\beta T$ is 40. A Gaussian kernel $G(\cdot)$ with a size of $3\times 3$ and a standard deviation $\sigma=0.5$ is used to smooth the attention map. To reduce the attention scores of non-maximum values, $\gamma$ is set to 0.85. We use spaCy [13] which is a library for advanced NLP, to extract noun tokens based on the part of speech of each word in the caption. All experiments are conducted on a NVIDIA A100 GPU.
Datasets.
As discussed in Section 4.1 of the main paper, we define Attend-Condition to further validate our method’s effectiveness in generating multi-object scenes under the paradigm of local control. In Table  3, we list the animals and objects inspired by [2].
Animals
cat, dog, bird, bear, horse, elephant, monkey,
frog, turtle, rabbit, mouse
Objects
glasses, crown, suitcase, chair, balloon,
car, bowl, bench, clock, apple
Table 3: Datasets. We list the animals and objects used in Attend-Condition dataset.
In the COCO [19] validation dataset, we use the largest segment region in each image as the local spatial information for inference. Specifically, we generate a binary mask from the image’s largest segment region. This mask is then combined with the global control condition, thereby defining the local control condition. The text prompt is randomly selected from one of the five captions corresponding to the current image. Then the local control condition and text prompt are fed into our pipeline.
8 More Ablation Studies
Control Concept Matching
The figure 8 shows the necessity of selecting concepts which match the local control region. The middle images correspond to manually specified $\text{C}^{t}_{\text{control}}$ of “cat” and “dog” , respectively. It can be observed that the content generated in the local region aligns with the manually specified objects. However, this naive method leads to the forced generation of objects that should not be in local region, affecting the model’s ability to align with the control condition. The results for both cat and dog generation are bad. The right image demonstrates the superiority of Control Concept Matching. It adaptively matches the concept during denoising process. More examples are shown in the Figure 9. Furthermore, Figure 10 shows the impact of different $\beta$ values on image quality.
Figure 8: Comparison of $\text{C}^{t}_{\text{control}}$ obtained by manually specifying and Control Concept Matching. Our method does not require manually specifying $\text{C}^{t}_{\text{control}}$ and produces higher quality images.
Focused Token Response
In Figure 11, we demonstrate the impact of different values of $\gamma$ on the synthesized image. When $\gamma$ is 1, it is difficult to generate images with both “cat” and “dog” , and alignment with the local condition is not achieved. When $\gamma$ is set to 0.85, the generated images meet expectations. As $\gamma$ decreases from 0.7 to lower values, the overall quality of the images significantly deteriorates.
9 More Visualization Results
We present visualization results under local control on both COCO validation and Attend-Condition datasets. Figure 14 displays Attend-Condition examples, showcasing the ability to align with captions and local conditions across various concept combinations. In Figure 13, selected original images of varying sizes from the COCO validation set and their corresponding generated images are displayed. Our method controls the generation of specific image areas based on user-defined local conditions, with the rest of the image following the original text prompt. Moreover, in the last two rows of Figure 13, even when the local control condition almost covers the entire image, our method still correctly generates images that comply with both the prompt and control condition. This demonstrates that our approach can generalize to global control tasks. Figure 12 provides additional visualization results applied under various conditions and prompts.
10 The Necessity of the Local Control
We emphasize the necessity of the local control paradigm. In Figure 15, when global control information at the image-level is input, the model can generate good content for the main subject as dictated by the control information. However, other objects described in the prompt depend on the control information outside the main subject’s control area. When the caption does not match well with the remaining control information, it can result in failure cases or produce somewhat strange content.
Figure 9: The first row represents manually specifying $\text{C}^{t}_{\text{control}}$ as “dog”, while the second row employs the proposed Control Concept Matching method. Our method generates concepts in the local region. At the same time, it correctly creates other parts of the image according to the prompt. It does not generate the same concept twice or create extra concepts.
Figure 10: Different $\beta$ values affect image quality. Our focus is on stabilizing the selection of $\text{C}^{t}_{\text{control}}$. With a higher $\beta$, emphasizing local region objects too early leads to poorer quality in these areas. Conversely, a lower $\beta$ means the objects assigned within the local region during the denoising process change, resulting in unstable content generation in the images. We have chosen 0.8 for our method.
Figure 11: The impact of different values of $\gamma$ on the synthesized image. At $\gamma=0.85$, image quality is optimal, and it declines below 0.7 and struggles with combination object generation at $\gamma=1$.
Figure 12: More visualization results. Our method achieves excellent results under a variety of conditions and prompts.
Figure 13: Synthetic samples with local conditions on COCO validation set. Our method accurately generates images under different control regions. Additionally, when the control covers almost the entire image (in the last two rows), our approach is still effective.
Figure 14: Synthetic samples with local conditions on Attend-Condition. The images showcase our method’s ability to align with captions and local conditions across diverse concept combinations.
Figure 15: Under global conditions, “cat” can be generated normally, while “toy vehicle” can only be generated based on background control conditions. The final generated result is unsatisfactory and does not meet the user’s needs. In the local control paradigm, users can specify “cat” to use the control condition, while the remaining areas are generated correctly based on the prompt.